OpenStudy (anonymous):

Consider the function f(x) = 8 √{ x} + 9 on the interval [ 2 , 8 ] (A) Find the average or mean slope of the function on this interval (B) By the Mean Value Theorem, we know there exists at least one c in the open interval ( 2 , 8 ) such that f′( c) is equal to this mean slope. Find all values of c that work and list them (separated by commas) in the box below.

OpenStudy (amistre64):

this one here?

OpenStudy (anonymous):

yea this is one of the ones i am having trouble on...

OpenStudy (amistre64):

mean slope means what? the slope of the line from end point to end point?

OpenStudy (anonymous):

i don't know i check all my notes and no where do i see mean slope so i don't know why the teacher assigned this problem

OpenStudy (amistre64):

I believe that it means the average slope between the intervals and is found by taking the slope of the line between the end points.

OpenStudy (amistre64):

then you find a point in the interval that touches the graph at one point with that slope

OpenStudy (anonymous):

so what would be the equation to find the slope between the intervals

OpenStudy (amistre64):

f(8) - f(2) change in y -------- over 8-2 change in x

OpenStudy (amistre64):

this is how you find any slope of a line :) the change in y over the change in x

OpenStudy (anonymous):

ok then what next?

OpenStudy (amistre64):

whats our mean slope? I get: 4sqrt(2) ------ 3

OpenStudy (amistre64):

when f'(c) = 4sqrt(2)/3 then we have found the values for "x" that satisfy the problem

OpenStudy (anonymous):

thats what i got...it was correct

OpenStudy (amistre64):

f(x) = 8x^(1/2) + 9 whats our derivative?

OpenStudy (amistre64):

4/sqrt(x) maybe?

OpenStudy (anonymous):

yes it is

OpenStudy (amistre64):

yay!!... 4sqrt(x) 4sqrt(2) ------ = -------- we need to solve this then x 3

OpenStudy (amistre64):

12 sqrt(x) = 4x sqrt(2) ; divide both sides by 4 3 sqrt(x) = x sqrt(2)

OpenStudy (anonymous):

the answer is 4.5....

OpenStudy (amistre64):

did you solve it, or is that what we are going for?

OpenStudy (anonymous):

i solved it...well after you help of course...i started to see the problem working out...i got another...

OpenStudy (amistre64):

I was gonna square both sides and do some magic :)

OpenStudy (anonymous):

\[f(x)=(x-3)(x-7)^3+8\] we are looking for absolute max and min for intervals at [1,4] and [1,8] and [4,9]

OpenStudy (amistre64):

.... i got 9/2...... so nyahh!!

OpenStudy (amistre64):

same function?

OpenStudy (amistre64):

.......ok, i can read..... sometimes :)

OpenStudy (amistre64):

its a quartic equation which will resemble a "W" perhaps

OpenStudy (amistre64):

just expand it and make it easier for ya...

OpenStudy (anonymous):

ok..lol... i got \[f'(x)=(x-7)^3+(x-3)(3)(x-7)^2\]

OpenStudy (amistre64):

f(x) = x^4 -24x^3 +210x^2 -784x +1037 f'(x) = 4x^3 -72x^2 +420x -784 :)

OpenStudy (amistre64):

your way might be easier to read :)

OpenStudy (anonymous):

ok so i would make the equation equal to zero right???

OpenStudy (amistre64):

yep..

OpenStudy (amistre64):

x = 7 is a zero, thats easy to see

OpenStudy (anonymous):

and would 3 be a zero

OpenStudy (amistre64):

3 would be a zero + (-7)^3

OpenStudy (amistre64):

take the second derivative to see how the graph is behaving

OpenStudy (amistre64):

f''(x) = 12x^2 -144x +420

OpenStudy (amistre64):

factor out a 12.. 12(x^2 - 12x + 35)

OpenStudy (amistre64):

12(x-7)(x-5) is the factored form

OpenStudy (anonymous):

ok and how would i use that to find the max and min between all three intervals

OpenStudy (amistre64):

draw a number line straight across the paper......

OpenStudy (amistre64):

<..................5......................7....................>

OpenStudy (amistre64):

at 5 and 7 we have inflection points...

OpenStudy (anonymous):

ok but our first interval were looking at is [1,4]

OpenStudy (amistre64):

<..................5......................7....................> --- +++ +++ --- --- +++ ---------------------------------- + - + when we multiply these "zones" together, we get these results

OpenStudy (amistre64):

do you undertand this diagram?

OpenStudy (anonymous):

yea...how do we apply the intervals to this...

OpenStudy (amistre64):

[1,4] is in an area that is concave up, so if anything it would have a MIN in it someplace

OpenStudy (amistre64):

Now, are we looking for the "highest and or lowest point in the intervals?

OpenStudy (anonymous):

i am going to break up the question like it is.... f(x)=(x−3)(x−7)^3+8 a. interval [1,4] absolute max: absolute min: this is how all of the other ones are

OpenStudy (amistre64):

our graph is acting like this:

OpenStudy (amistre64):

solve for f(1) and f(4) to see what our values are.

OpenStudy (amistre64):

f(1) = 440 if I did it right

OpenStudy (amistre64):

f(4) = -19

OpenStudy (amistre64):

the max is at (1,440)

OpenStudy (amistre64):

here a thought, plug in 4 into the derivative and see if we get a zero :)

OpenStudy (amistre64):

yep, 4 is a zero for f'(x)...im a genuis :)

OpenStudy (amistre64):

[1,4] (1 ,440) = max and (4,-19) = min you agree?

OpenStudy (amistre64):

and the graph doesnt have an more "bumps in it, just some transitional bends.....

OpenStudy (anonymous):

yea the next interval is the same and the last one the max is (9,56) and min is (4,-19)

OpenStudy (anonymous):

\[f(x)=(3+9\ln (x))/x\] i need to find the x coordinate of the absolute max...

OpenStudy (amistre64):

start a new question, I think this ones lagging alot becasue of all the "replies" its trying to keep track of :)